# Arnold-Thom conjecture for the arrival time of surfaces

### v1

@@ 1. Introduction
In [T89], Thom conjectured that if a gradient flow line for an analytic function has a limit point, then the secants of this gradient flow line converge. Later, Arnold [A04] proposed a stronger conjecture suggesting the convergence of its tangent directions. This is now known as the Arnold-Thom gradient conjecture.
Conjecture 1.1 (Arnold-Thom).
@@
Thom’s conjecture is based on the Łojasiewicz theorem about the zero-order asymptotic behavior of gradient flow lines. In [L84], Łojasiewicz proved that if a gradient flow line of an analytic function has a limit point, then the flow line has finite length, and hence has a unique limit point. His proof used two important inequalities for analytic functions, now known as the Łojasiewicz inequalities [L]. Based on the work of Łojasiewicz, Kurdyka-Mostowski-Parusiński proved Thom’s conjecture in [KMP] by showing that the curve of secants has finite length after being projected to the unit sphere. However, Conjecture 1.1, carrying finer first-order asymptotic information of gradient flow lines, is still open.
In this paper, we study the Łojasiewicz theorem and the Arnold-Thom conjecture for solutions to the arrival time equation
@@
on mean convex domains in $\mathbb{R}^{n+1}.$ This is a degenerate elliptic partial differential equation, so its solutions are far from being analytic. By the works of Evans-Spruck [ES] (cf. [CGG]) and Colding-Minicozzi [CM16], this equation admits a unique $C^{1,1}$ solution in any smooth compact mean convex domain given a Dirichlet boundary condition. This equation is connected to mean curvature flows in the following geometric way. If $u$ is an arrival time function in a mean convex domain $\Omega\subseteq\mathbb{R}^{n+1},$ that is, $u$ solves the equation (1.2) in $\Omega,$ then its level set $M_{t}:=\left\{x\in\Omega:u(x)=t\right\}$ evolves by the mean curvature flow. In other words, the mean curvature vector $\textbf{H}=-H\mathbf{n}$ of $M_{t}$ satisfies $\partial_{t}x=\textbf{H}.$ Here, a smooth domain is mean convex if its boundary has non-negative mean curvature with respect to its outer unit normal.
@@ 1.1. Main results
Let us consider gradient flow lines for the arrival time function $u$, which is a solution to (1.2). In this context, we prove the Łojasiewicz theorem and solve the Arnold-Thom conjecture for arrival time functions in $\mathbb{R}^{3}$ (cf. [CM19, Conjecture 1.4]).
@@
This theorem is a consequence of a more general result that we show in $\mathbb{R}^{n+1}$ with $n\geq 2$. In fact, we prove the Łojasiewicz theorem and the Arnold-Thom conjecture when a limit point of a gradient flow is a spherical or neck singularity. The precise statement is recorded in Theorem 3.11. The condition of Theorem 3.11 is automatically true if the initial closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ satisfies one of the following conditions (cf. [HS09, CM162, HK2, G24]).
@@
Thus, we prove the Łojasiewicz theorem and solve the Arnold-Thom conjecture in $\mathbb{R}^{n+1}$ in the two-convex case and the low-entropy case.
Another situation that we can work on in the higher dimensional case is the non-degenerate condition, which is also exploited in [SS]. A cylindrical singularity is called non-degenerate if, roughly speaking, the flow looks like a cylinder when zoomed in in any directions and scales. That is, all limit flows [^1] In the study of geometric flows, limiting models are described using the notion of limit flows and tangent flows. See Definition 2.3. are cylinders. A priori, this assumption does not exclude the possibility that under different scales and directions, the flow looks like cylinders of different sizes or axes. We prove that this cannot happen. Using this result, we then prove the Łojasiewicz theorem and the Arnold-Thom conjecture for gradient flow lines having a non-degenerate cylindrical singularity as a limit point.
@@
Let $M_{t}$ be a closed mean convex mean curvature flow in $\mathbb{R}^{n}.$ If $\gamma$ is a gradient flow line for the arrival time function of $M_{t}$ with a non-degenerate limit point, then the limit flow is unique, the limit point is unique, and the unit tangents of $\gamma$ converge.
Examples of non-degenerate cylindrical singularities were constructed [GS09]. The stability results of Schulze-Sesum [SS] implies that there are more. A related notion of non-degeneracy was discussed in [SX4].
Finally, we deal with general mean curvature flows with neck singularities when mean curvatures change signs or even when their level set flows fatten. The key is the mean convex neighborhood theorem proven by Choi-Haslhofer-Hershkovits-White [CHHW] (cf. [CHH]), which implies that flows behave nicely near neck singularities. This means that no matter how strangely a flow line goes, it will be constrained by the ordinary arrival time equation as in the mean convex case when the flow line approaches a neck singularity. Therefore, we can apply the arguments in the mean convex case to get the convergence of tangents in this general setting. In particular, with the recent breakthrough of Bamler-Kleiner [BK23] about the multiplicity one conjecture, the Łojasiewicz theorem and the Arnold-Thom conjecture are true for a mean curvature flow starting from a two-sphere (based on [W95, Br16, CHH]) or a generic surface in $\mathbb{R}^{3}$ (based on [CCMS, CCS, SX1, SX2]).
@@
As we will see later, solutions to the arrival time equation (1.2) may not be $C^{2},$ let alone analytic. Without mean convexity, solutions to the level set equations are even known to be only Lipschitz continuous. Thus, the results above support the principle “that solutions of degenerate equations behave as though they are analytic even when they are not” proposed by Colding-Minicozzi [CM19].
1.2. Previous works and proof ideas
@@
We mention some earlier works. The method of using level sets to study geometric flows started from Osher-Sethian [OS88], Evans-Spruck [ES], Chen-Giga-Goto [CGG], and Ilmanen [I94]. In the case of mean convex mean curvature flows, Evans-Spruck further showed that level set flows can be given by functions with a separation of variables, leading to the arrival time equation (1.2). They proved that there is always a Lipschitz viscosity weak solution to this degenerate elliptic partial differential equation by approximating it using solutions to perturbed elliptic equations.
@@
The optimal regularity of solutions to (1.2) has been a subtle issue for a long time. When the singularity is spherical, based on the works of Huisken [H93], Sesum [S08], and Strehlke [S20], we know that the optimal regularity is $C^{2}$ when the dimension of flows is at least two. [^2] For the curve case, the solutions to (1.2) are known to be at least $C^{3}$ based on the result of Kohn-Serfaty [KS06]. This turns out to be false when the singularity is not spherical. In fact, Ilmanen [I92] proved that even for a typical dumbbell example constructed by Grayson [G89], its arrival time function is not $C^{2}$ at the neck singularity. In [CM16, CM18], Colding-Minicozzi improved the result and proved that solutions to (1.2) are always $C^{1,1}$ and twice differentiable. That is, for an arrival time function $u,$ its Hessian exists everywhere, though it may not be continuous. Moreover, at a cylindrical singularity $p$ modeled on $S^{n-k}\left(\sqrt{2(n-k)}\right)\times\mathbb{R}^{k}$ of a mean convex flow $M_{t},$ its arrival time function is $C^{2}$ at $p$ if and only if the singular set near the point is a $k$-dimensional $C^{1}$ submanifold (cf. Theorem 2.4 and [G24]). In both spherical and cylindrical cases, Sun-Xue [SX4, SX3] showed that these low-regularity phenomena are generic. There are only few examples with $C^{2}$ regularity, for instance, the $O(k+1)\times O(n-k+1)$ symmetric marriage ring $S^{k}\times S^{n-k}.$
@@
For a $C^{2}$ arrival time function, in [CM19], Colding-Minicozzi proved that its gradient flow lines satisfy the Arnold-Thom conjecture. In this delicate work, they first proved an optimal Łojasiewicz inequality for a $C^{2}$ arrival time function near its singularity. This is already a non-trivial work, implying the Łojasiewicz theorem about the uniqueness of limits points for gradient flow lines. Based on this, they reduced the Arnold-Thom conjecture to a sharp convergence rate for rescaled mean curvature flows converging to a cylinder. Obtaining such a sharp convergence rate is technical, requiring a careful analysis of the kernel of the stability operator (cf. [CM182]).
@@
In general, solutions without $C^{2}$ regularity near cylindrical singularities are more prevalent. Therefore, we need to pursue new methods. We mention the proof ideas and how our results are related to previous works. The first subtle point is that when the arrival time function $u$ is not $C^{2},$ a priori, we don’t even know the uniqueness of limit points, that is, the Łojasiewicz theorem, as there is no Łojasiewicz inequality for $u.$ In fact, if the uniqueness is not known, it does not make sense to formulate Thom’s original conjecture. We obtain this uniqueness result by applying the structure theorems for singularities established by Colding-Minicozzi [CM162, CM18]. This method uses special properties of one-dimensional curves and does not extend to singularities modeled on cylinders with more than one Euclidean factors.
@@
Our theorems about the Arnold-Thom conjecture are based on two important ingredients. The first one is the regularity of arrival time functions [CM18, CM19]. Given a mean convex flow $M_{t}$ having a neck singularity $p,$ to resolve the Arnold-Thom conjecture, we need to deal with the case when the arrival time function $u$ is not $C^{2}$ near $p.$ In this case, although there is no optimal regularity, using the structural result of singular sets [CM162], Colding-Minicozzi [CM18] were able to show that the function $u$ still behaves nicely when approached away from the kernel of the Hessian of $u,$ that is, the axis of the tangent cylinder. Thus, if a gradient flow line approaches a neck singularity transversely, one would expect the asymptotic behavior to be still nice. This will be related to finer study near a neck singularity, which will be addressed next.
@@
The second ingredient is the thorough understanding of the flow behavior near neck singularities. By the works of Brendle-Choi [BC19, BC21], Angenent-Daskalopoulos-Sesum [ADS20] and Choi-Haslhofer-Hershkovits-White [CHHW], the canonical neighborhoods near a neck singularity are fully classified: along a gradient flow line, a limit flow can only be the shrinking cylinder or the translating bowl soliton constructed in [AW] (cf. [W03, W11, H15]) up to rotations and translations. If limit flows are all shrinking cylinders along a gradient curve, then we can argue that the gradient flow line approaches the neck singularity transversely, so the situation is reduced to the work of Colding-Minicozzi. Thus, the task becomes classifying possible limit flows along a fixed gradient flow line. What we finally obtain is a uniqueness result of limit flows. That is, along a fixed gradient flow line, if one limit flow is a shrinking cylinder, then all the others must be. Once we have this uniqueness result, we can further prove the corresponding uniqueness when the limit flows are translating bowl solitons. In this case, the flow approaches the singularity slowly and non-transversely, so we have to deal with it separately.
@@
We remark that there are many examples of neck singularities at which their arrival time functions are not $C^{2}.$ As we mention, Sun-Xue [SX4] proved that this is a generic phenomenon (cf. [SS]), and there are many explicit constructions [AV, AAG]. There are also works about infinite dimensional versions of Thom’s conjecture, see [CH23].
@@
We organize the paper as follows. Section 2 provides a preliminary introduction to mean convex mean curvature flows. In Section 3, we first show the uniqueness of limit points for a gradient flow line with a neck limit point, proving the Łojasiewicz theorem. We then prove the uniqueness of limit flows along such a gradient flow line and use it to prove the Arnold-Thom conjecture for neck singularities, which implies Theorem 1.3. In Section 4, we deal with general cylindrical singularities that are non-degenerate and prove Theorem 1.4. Section 5 concludes with the results about general mean curvature flows with neck singularities, including flows starting from two-spheres and generic closed hypersurfaces.
Acknowledgement
@@
The authors are grateful to Prof. Bill Minicozzi for several insightful comments. They would like to thank Ao Sun and Zhihan Wang for helpful conversations. During the project, Lee was partially supported by NSF Grant DMS 2005345.
@@ 2.1. Mean convex flows
In the work of Chen-Giga-Goto [CGG] and Evans-Spruck [ES], they showed the existence and uniqueness of viscosity solutions to the level set flow equation
@@
in a compact set in $\mathbb{R}^{n+1}.$ When a given hypersurface $M$ is mean convex, Evans-Spruck [ES] further showed that the solution to (2.1) can be given by a separation of variables, that is, of the form $v(x,t)=u(x)-t.$ Thus, the new function $u$ in this context satisfies the arrival time equation (1.2), and the mean convex level set flow $M_{t}$ starting from $M$ can be given by the level sets of $u,$ which is defined on the smooth closed mean convex domain $\Omega$ with $\partial\Omega=M.$ If we let $t=-T$ be the starting time, since $M_{-T}=M$ is closed and the flow is in a Euclidean space, all singularities occur in finite time. Thus, the arrival time function $u$ constructed above has a finite supremum. That is, there exists $T_{\rm large}>0$ such that $M_{t}=\emptyset$ for all $t\geq T_{\rm large}.$ Thus, for a closed mean convex level set flow, we will always assume the time interval to be $[-T,\infty).$
@@
(1) (Neck singularity) A singularity $(x_{0},t_{0})$ of the flow is called a cylindrical singularity if the tangent flow at $(x_{0},t_{0})$ is modeled on a multiplicity one $k$-cylinder. When $k=1,$ it is called a neck singularity.
@@
Note that for a tangent flow modeled on a $k$-cylinder, it represents a singularity only when $k<n.$ Also, cylindrical tangent flows are unique by [CM15]. That is, if a tangent flow at a singularity is given by one generalized cylinder, then any other tangent flow at the singularity is given by the same cylinder (with the same axis). Thus, the definition above makes sense.

### v2

# Arnold–Thom conjecture for the arrival time of surfaces

@@ 1. Introduction
In [T89], Thom conjectured that if a gradient flow line for an analytic function has a limit point, then the secants of this gradient flow line converge. Later, Arnold [A04] proposed a stronger conjecture suggesting the convergence of its tangent directions. This is now known as the Arnold–Thom gradient conjecture.
Conjecture 1.1 (Arnold–Thom).
@@
Thom’s conjecture is based on the Łojasiewicz theorem about the zero-order asymptotic behavior of gradient flow lines. In [L84], Łojasiewicz proved that if a gradient flow line of an analytic function has a limit point, then the flow line has finite length, and hence has a unique limit point. His proof used two important inequalities for analytic functions, now known as the Łojasiewicz inequalities [L]. Based on the work of Łojasiewicz, Kurdyka–Mostowski–Parusiński proved Thom’s conjecture in [KMP] by showing that the curve of secants has finite length after being projected to the unit sphere. However, Conjecture 1.1, carrying finer first-order asymptotic information of gradient flow lines, is still open.
In this paper, we study the Łojasiewicz theorem and the Arnold–Thom conjecture for solutions to the arrival time equation
@@
on mean convex domains in $\mathbb{R}^{n+1}.$ This is a degenerate elliptic partial differential equation, so its solutions are far from being analytic. By the works of Evans–Spruck [ES] (cf. [CGG]) and Colding–Minicozzi [CM16], this equation admits a unique $C^{1,1}$ solution in any smooth compact mean convex domain given a Dirichlet boundary condition. This equation is connected to mean curvature flows in the following geometric way. If $u$ is an arrival time function in a mean convex domain $\Omega\subseteq\mathbb{R}^{n+1},$ that is, $u$ solves the equation (1.2) in $\Omega,$ then its level set $M_{t}:=\left\{x\in\Omega:u(x)=t\right\}$ evolves by the mean curvature flow. In other words, the mean curvature vector $\textbf{H}=-H\mathbf{n}$ of $M_{t}$ satisfies $\partial_{t}x=\textbf{H}.$ Here, a smooth domain is mean convex if its boundary has non-negative mean curvature with respect to its outer unit normal.
@@ 1.1. Main results
Consider a gradient flow line for the arrival time function $u$, which is a solution to (1.2). In this context, we prove the Łojasiewicz theorem and solve the Arnold–Thom conjecture for arrival time functions in $\mathbb{R}^{3}$ (cf. [CM19, Conjecture 1.4]).
@@
This theorem is a consequence of a more general result that we show in $\mathbb{R}^{n+1}$ with $n\geq 2$. In fact, we prove the Łojasiewicz theorem and the Arnold–Thom conjecture when a limit point of a gradient flow is a neck singularity. The precise statement is recorded in Theorem 3.16. The condition of Theorem 3.16 is automatically true if the initial closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ satisfies one of the following conditions (cf. [HS09, CM162, HK2, G24]).
@@
Thus, we prove the Łojasiewicz theorem and solve the Arnold–Thom conjecture in $\mathbb{R}^{n+1}$ in the two-convex case and the low-entropy case.
Another situation that we can work on in the higher dimensional case is the non-degenerate condition, which is also exploited in [SS]. A cylindrical singularity is called non-degenerate if, roughly speaking, the flow looks like a cylinder when zoomed in in any directions and scales. That is, all limit flows [^1] In the study of geometric flows, limiting models are described using the notion of limit flows and tangent flows. See Definition 2.3. are cylinders. A priori, this assumption does not exclude the possibility that under different scales and directions, the flow looks like cylinders of different sizes or axes. We prove that this cannot happen. Using this result, we then prove the Łojasiewicz theorem and the Arnold–Thom conjecture for gradient flow lines having a non-degenerate cylindrical singularity as a limit point.
@@
Let $M_{t}$ be a closed mean convex mean curvature flow in $\mathbb{R}^{n+1}.$ If $\gamma$ is a gradient flow line for the arrival time function of $M_{t}$ with a non-degenerate limit point, then the limit flow is unique, the limit point is unique, and the unit tangents of $\gamma$ converge.
Examples of non-degenerate cylindrical singularities were constructed [GS09]. The stability result of Schulze–Sesum [SS] implies that there are more. A related notion of non-degeneracy was discussed in [SX4, SWX].
Finally, we deal with general mean curvature flows with neck singularities when mean curvatures change signs or even when their level set flows fatten. The key is the mean convex neighborhood theorem proven by Choi–Haslhofer–Hershkovits–White [CHHW] (cf. [CHH]), which implies that flows behave nicely near neck singularities. This means that no matter how strangely a flow line goes, it will be constrained by the ordinary arrival time equation as in the mean convex case when the flow line approaches a neck singularity. Therefore, we can apply the arguments in the mean convex case to get the convergence of tangents in this general setting. In particular, with the recent breakthrough of Bamler–Kleiner [BK23] about the multiplicity one conjecture, the Łojasiewicz theorem and the Arnold–Thom conjecture are true for a mean curvature flow starting from a two-sphere (based on [W95, Br16, HW, CHH]) or a generic surface in $\mathbb{R}^{3}$ (based on [CCMS, CCS, SX1, SX2]).
@@
As we will see later, solutions to the arrival time equation (1.2) may not be $C^{2},$ let alone analytic. Without mean convexity, solutions to the level set equations are even known to be only Lipschitz continuous. Thus, the results above support the principle “that solutions of degenerate equations behave as though they are analytic even when they are not” proposed by Colding–Minicozzi [CM19].
1.2. Previous results and proof ideas
@@
We mention some earlier works. The method of using level sets to study geometric flows started from Osher–Sethian [OS88], Evans–Spruck [ES], Chen–Giga–Goto [CGG], and Ilmanen [I94]. In the case of mean convex mean curvature flows, Evans–Spruck further showed that level set flows can be given by functions with a separation of variables, leading to the arrival time equation (1.2). They proved that there is always a Lipschitz viscosity weak solution to this degenerate elliptic partial differential equation by approximating it using solutions to perturbed elliptic equations.
@@
The optimal regularity of solutions to (1.2) has been a subtle issue for a long time. When the singularity is spherical, based on the works of Huisken [H93], Sesum [S08], and Strehlke [S20], we know that the optimal regularity is $C^{2}$ when the dimension of flows is at least two. [^2] For the curve case, the solutions to (1.2) are known to be at least $C^{3}$ based on the result of Kohn–Serfaty [KS06]. This turns out to be false when the singularity is not spherical. In fact, Ilmanen [I92] proved that even for a typical dumbbell example constructed by Grayson [G89], its arrival time function is not $C^{2}$ at the neck singularity. In [CM16, CM18], Colding–Minicozzi improved the result and proved that solutions to (1.2) are always $C^{1,1}$ and twice differentiable. That is, for an arrival time function $u,$ its Hessian exists everywhere, though it may not be continuous. Moreover, at a cylindrical singularity $p$ modeled on $S^{n-k}\left(\sqrt{2(n-k)}\right)\times\mathbb{R}^{k}$ of a mean convex flow $M_{t},$ its arrival time function is $C^{2}$ at $p$ if and only if the singular set near the point is a $k$-dimensional $C^{1}$ submanifold (cf. Theorem 2.4 and [G24]). In both spherical and cylindrical cases, Sun–Xue [SX4, SX3] showed that these low-regularity phenomena are generic. There are only few examples with $C^{2}$ regularity, for instance, the $O(k+1)\times O(n-k+1)$ symmetric marriage ring $S^{k}\times S^{n-k}.$
@@
For a $C^{2}$ arrival time function, in [CM19], Colding–Minicozzi proved that its gradient flow lines satisfy the Arnold–Thom conjecture. In this delicate work, they first proved an optimal Łojasiewicz inequality for a $C^{2}$ arrival time function near its singularity. This is already a non-trivial result, implying the Łojasiewicz theorem about the uniqueness of limits points for gradient flow lines. Based on this, they reduced the Arnold–Thom conjecture to a sharp convergence rate for rescaled mean curvature flows converging to a cylinder. Obtaining such a sharp convergence rate is technical, requiring a careful analysis of the kernel of the stability operator (cf. [CM182]).
@@
In general, solutions without $C^{2}$ regularity near cylindrical singularities are more prevalent. Therefore, we need to pursue new methods. We mention the proof ideas and how our results are related to previous results. The first subtle point is that when the arrival time function $u$ is not $C^{2},$ a priori, we don’t even know the uniqueness of limit points, that is, the Łojasiewicz theorem, as there is no Łojasiewicz inequality for $u.$ In fact, if the uniqueness is not known, it does not make sense to formulate Thom’s original conjecture. We obtain this uniqueness result by applying the structure theorems for singularities established by Colding–Minicozzi [CM162, CM18]. This method uses special properties of one-dimensional curves and does not extend to singularities modeled on cylinders with more than one Euclidean factor.
@@
Our theorems about the Arnold–Thom conjecture are based on two important ingredients. The first one is the regularity of arrival time functions [CM18, CM19]. Given a mean convex flow $M_{t}$ having a neck singularity $p,$ to resolve the Arnold–Thom conjecture, we need to deal with the case when the arrival time function $u$ is not $C^{2}$ near $p.$ In this case, although there is no optimal regularity, using the structural result of singular sets [CM162], Colding–Minicozzi [CM18] were able to show that the function $u$ still behaves nicely when approached away from the kernel of the Hessian of $u,$ that is, the axis of the tangent cylinder. Thus, if a gradient flow line approaches a neck singularity transversely, one would expect the asymptotic behavior to be still nice. This will be related to finer study near a neck singularity, which will be addressed next.
@@
The second ingredient is the thorough understanding of the flow behavior near neck singularities. By the works of Brendle–Choi [BC19, BC21], Angenent–Daskalopoulos–Sesum [ADS20] and Choi–Haslhofer–Hershkovits–White [CHHW], the canonical neighborhoods near a neck singularity are fully classified: along a gradient flow line, a limit flow can only be the shrinking cylinder or the translating bowl soliton constructed in [AW] (cf. [W03, W11, H15]) up to rotations and translations. If limit flows are all shrinking cylinders along a gradient curve, then we can argue that the gradient flow line approaches the neck singularity transversely, so the situation is reduced to the work of Colding–Minicozzi. Thus, the task becomes classifying possible limit flows along a fixed gradient flow line. What we finally obtain is a uniqueness result of limit flows. That is, along a fixed gradient flow line, if one limit flow is a shrinking cylinder, then all the others must be. Once we have this uniqueness result, we can further prove the corresponding uniqueness when the limit flows are translating bowl solitons. In this case, the flow approaches the singularity slowly and non-transversely, so we have to deal with it separately.
@@
We remark that there are many examples of neck singularities at which their arrival time functions are not $C^{2}.$ As we mention, Sun–Xue [SX4] proved that this is a generic phenomenon (cf. [SS]), and there are many explicit constructions [AV, AAG]. There are also works about infinite dimensional versions of Thom’s conjecture, see [CH23].
@@
We organize the paper as follows. Section 2 provides a preliminary introduction to mean convex mean curvature flows. In Section 3, we first show the uniqueness of limit points for a gradient flow line with a neck limit point, proving the Łojasiewicz theorem. We then prove the uniqueness of limit flows along such a gradient flow line and use it to prove the Arnold–Thom conjecture for neck singularities, which implies Theorem 1.3. In Section 4, we deal with general cylindrical singularities that are non-degenerate and prove Theorem 1.4. Section 5 concludes with the results about general mean curvature flows with neck singularities, including flows starting from two-spheres and generic closed hypersurfaces.
Acknowledgment
@@
The authors are grateful to Prof. Bill Minicozzi for several insightful comments. They would like to thank Ao Sun and Zhihan Wang for helpful conversations and thank the anonymous referees for constructive comments and suggestions, which have helped improve this paper. During the project, Lee was partially supported by NSF Grant DMS 2005345.
@@ 2.1. Mean convex flows
In the work of Chen–Giga–Goto [CGG] and Evans–Spruck [ES], they showed the existence and uniqueness of viscosity solutions to the level set flow equation
@@
in a compact set in $\mathbb{R}^{n+1}.$ When a given hypersurface $M$ is mean convex, Evans–Spruck [ES] further showed that the solution to (2.1) can be given by a separation of variables, that is, of the form $v(x,t)=u(x)-t.$ Thus, the new function $u$ in this context satisfies the arrival time equation (1.2), and the mean convex level set flow $M_{t}$ starting from $M$ can be given by the level sets of $u,$ which is defined on the smooth closed mean convex domain $\Omega$ with $\partial\Omega=M.$ If we let $t=-T$ be the starting time, since $M_{-T}=M$ is closed and the flow is in a Euclidean space, all singularities occur in finite time. Thus, the arrival time function $u$ constructed above has a finite supremum. That is, there exists $T_{\rm large}>0$ such that $M_{t}=\emptyset$ for all $t\geq T_{\rm large}.$ Thus, for a closed mean convex level set flow, we will always assume the time interval to be $[-T,\infty).$
@@
(1) (Neck singularity) A singularity $(x_{0},t_{0})$ of the flow is called a cylindrical singularity if the tangent flow at $(x_{0},t_{0})$ is a multiplicity one $k$-cylinder. When $k=1,$ it is called a neck singularity.
@@
Cylindrical tangent flows are unique by [CM15]. That is, if a tangent flow at a singularity is given by one generalized cylinder, then any other tangent flow at the singularity is given by the same cylinder (with the same axis). Thus, the definition above makes sense.